Write an equation in standard form for the line perpendicular to 8x+y=5 that passes through the point (3, 4)
Answer by mananth(16946)

8 x + 1 y = 5
Find the slope of this line

1 y = -8 x + 5
Divide by 1
y = -8 x + 5
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -8

The slope of a line perpendicular to the above line will be the negative reciprocal 1/8
m1*m2=-1
The slope of the required line will be 1/8

m= 1/8 ,point ( 3 , 4 )
Find b by plugging the values of m & the point in
y=mx+b
4 = 3/ 8 + b
b= 29/8
m= 1/8
The required equation is y = 1/ 8 x +29/ 8
m.ananth@hotmail.ca